# Sequential Attention-based Sampling for Histopathological Analysis

### v2

@@ A.3 Additional Control Experiments
A.3.1 Alternative Models under an Observation Budget
We compared the performance of SASHA with competing sota models like ACMIL [35] and DTFD [34] under the same observation budget of either 10% or 20%. That is, we predicted slide labels by having these sota models examine only a limited fraction of patches in each WSI at high resolution, matching the fraction of the WSI sampled by SASHA.
Each model was trained with a combination of low-resolution and high-resolution patch features and tested, similarly, with a combination of high and low resolution patch features in the same proportion as of training. Specifically, for a given slide, if $N$ is the total number of low-resolution patches and each low-resolution patch corresponds to $k$ high-resolution patches, then for that slide, the input comprised of a total of $N_{L}+N_{H}$ patches comprising $N_{L}=\left(N-\left\lceil f\cdot N\right\rceil\right)$ low-resolution and $N_{H}=\left\lceil f\cdot N\right\rceil\cdot k$ high-resolution patches, where $f\in\{0.1,0.2\}$. Because low and high-resolution patches were of the same pixel dimensionality $(256,256,3)$, the pretrained ViT model (Section 3) was used to extract features from each such patch; this yielded a feature vector $W\in\mathbb{R}^{(N_{L}+N_{H})\times d}$ which was provided as input to each model for label prediction, both during training and evaluation.
@@ A.3.1 Alternative Models under an Observation Budget
ACMIL - 0.1 0.841$\pm$0.036 0.828$\pm$0.034 0.773$\pm$0.038
DTFD - 0.1 0.829$\pm$0.013 0.835$\pm$0.039 0.829$\pm$0.013
SASHA - 0.1 0.901$\pm$0.021 0.918$\pm$0.014 0.856$\pm$0.031
ACMIL - 0.2 0.838$\pm$0.035 0.857$\pm$0.030 0.773$\pm$0.043
DTFD - 0.2 0.862$\pm$0.027 0.878$\pm$0.030 0.862$\pm$0.027
SASHA - 0.2 0.953$\pm$0.017 0.979$\pm$0.008 0.937$\pm$0.024
@@
As shown in Table 7, such a strategy of sparsely sampling patches substantially reduced the performance of ACMIL and DTFD models. In each case, SASHA sampling either 10% (0.1 fraction) or 20% (0.2 fraction) of the patches at high resolution outperformed the sota models with the corresponding, identical observation budget. In other words, competing sota models failed to live up to SASHA’s performance under a comparable observation budget for high-resolution patches.
A.3.2 Alternative Action Policies during Inference
In the current SASHA implementation, the inference strategy follows a deterministic, greedy policy wherein, for a given state, the action with the highest probability is always selected. This means the agent always picks the most likely action as per the learned policy distribution during inference (but not during training, see Section A.2.3).
@@ A.3.2 Alternative Action Policies during Inference
Comparison of different sampling strategies is shown in Table 8. Interestingly, the default deterministic policy of SASHA – a greedy strategy of selecting the action with the highest probability – outperformed all other stochastic policies. This behavior was observed empirically, where the agent consistently selects action with highest probability given current state $S_{t}$, resulting in more informative state transitions from $S_{t}$ to $S_{t+1}$ and improved classification performance, as shown in Table 8 . In addition, the stochastic policy that had the greatest degree of freedom for sampling (Stochastic - full distribution) outperformed, by and large, the other policies (top-$k$ and top-$p$); this was understandable because the latter policies impose more constraints on the action space for sampling.
@@
A.3.3 Alternative Reward Schedules during Training
The default SASHA model employs a step-wise reward during training, where the agent receives a reward at every time step of the episode associated with the label prediction loss at that time step (see Section A.2.3). In contrast, a previous competing model (RLogist [36]) employed a terminal reward strategy, providing feedback only at the end of each episode.
We tested this alternative reward schedule employing only a terminal reward at the end of each episode, both during training and inference. As shown in Table 9, the RL agent trained with reward at each time step significantly outperformed an agent trained using terminal rewards. We propose that this occurs because our reward formulation – which is designed to minimize the classification loss at each time step – encourages the model to sample more tumor-relevant patches during inference. Confirming this prediction SASHA’s default reward schedule sampled, on average, more patches with a high attention (relevance) score, as compared to the terminal reward schedule (Table 9).
Reward Schedule Accuracy AUC F1 Top-k attention score patches $\cap$ RL actions Avg attention score of RL actions
@@ A.3.3 Alternative Reward Schedules during Training
Table 9: Comparison of a terminal reward schedule (row 1) with SASHA’s default schedule of reward at each time step (row 2), during training (see text for details). Column 5 tells the average patches per slide that are selected by RL agent are in top-k according to the attention score given by HAFED model. Column 6 tells the average attention score of the selected patches by RL agent. (see Fig. 3f-g for details). Values denote the performance of the SASHA-0.2 model on the CAMELYON16 dataset. Bold: the best performing reward schedule.
A.3.4 Analysis of RL Agent Behavior
Figure 4: Evaluation of model performance on tumor slides from the test set, categorized into four groups. The first three columns correspond to correctly classified tumor slides (TS) by both SASHA and HAFED across increasing tumor fraction (TF) ranges, while the fourth column represents TS misclassified by SASHA but correctly predicted by HAFED. The x-axis denotes the visited patch fraction, binned with a width of 0.005. For each slide, we compute the average entropy, average tumor class prediction probability, and maximum cumulative tumor patch hit ratio (Eq. 1) within each bin. These metrics are then averaged across all slides, with solid lines indicating the mean and shaded regions representing the standard deviation.
@@ A.3.4 Analysis of RL Agent Behavior
First, we asked: how quickly does the RL agent visit tumor patches, and how does the model’s confidence evolve over time steps in an episode? We first performed this analysis for slides that were correctly classified by SASHA, separately for 3 classes of WSIs with varying tumor fractions (TF, Fig. 4, columns 1-3). For this, we defined a cumulative tumor patch Hit Ratio using
Hit $\text{Hit Ratio}(t)=\frac{\text{Number of tumor patches sampled up to time}\ t}{\min\left(\text{Observation Budget},\ \text{Total Tumor Patches in WSI}\right)}.$ (1)
@@
For both variants of the model (SASHA-0.1 and SASHA-0.2), on the slides with very sparse tumor regions (TF<0.1), the classifier’s output favoring the tumor class reaches a high value ($\sim$0.8-1.0) even before it vists 5% of all patches at high-resolution (Fig. 4, column 1, green trace). In addition, its predictive entropy drops progressively as more patches are sampled (Fig. 4, column 1, green trace). Surprisingly, even though the peak of the classifier output is reached earlier for SASHA-0.1 as compared to the SASHA-0.2 model, a longer observation budget produces higher accuracies, as well as better calibrated predictions, as shown in the main text Table 1 and Fig. 3-h.
As the budget increases from 10% to 20%, the tumor hit ratio improves correspondingly. With a 10% budget, SASHA tends to misclassify slides with small tumor regions (see F1 score in Table 1), highlighting the importance of increasing the sampling budget. At 20%, SASHA correctly classifies nearly all slides that were correctly predicted by HAFED, demonstrating improved robustness across varying tumor fractions.
As the fraction of tumor patches increases (0.1<TF<0.6 and TF>0.6), the model output favoring the tumor class reaches peak upon visiting progressively miniscule fractions of high-resolution patches (<0.025% or 0.01%, Fig. 4, columns 2-3, green traces). The predictive entropy also goes to near zero at these fractions indicating confident model predictions with very few samples.
For Fig. 4, column 2—where the tumor fraction (TF) lies in the range $0.1\leq\text{TF}<0.6$—the tumor hit ratio increases slightly as the budget increases. However, the improvement is minimal because most tumor patches are already captured within the initial 10% of patches selected by the RL agent. This suggests that for WSIs with moderate tumor presence, visiting just 10% of the patches is often sufficient.
@@
We illustrate this with an example slide Fig. 5. For this slide, the RL agent successfully identifies tumor patches, with the first tumor patch being visited at a fraction of $0.0085$ and the last at $0.03$. For the entire cohort of such slides tumor patch visits ranged from 0.001-0.099 fraction for SASHA-0.1 and 0.009-0.034 fraction for SASHA-0.2 (Fig. 4, column 4, orange trace). In contrast, when patches are randomly sampled for the same slide, the first tumor patch is typically hit after visiting $0.1412\pm 0.121$ fraction of the patches. This demonstrates that the RL agent is able to localize tumor regions significantly earlier than random sampling.
@@
A.4 Computation of WSI compressability, Inference time and ECE
Whole Slide Images (WSIs) are typically giga-resolution, requiring substantial storage space. For laboratories or hospitals maintaining long-term archives, this poses a significant storage challenge. However, it is not necessary to store WSIs in raw image format. By storing them as extracted feature representations, a considerable reduction in storage requirements can be achieved.
To quantify the data reduction achieved during WSI processing, we define the WSI compressibility factor as the ratio between the original image size at full magnification and the size of the final feature representation used for classification. Let the raw image size be $W\times H\times 3$, representing the width, height, and RGB color channels.
Models such as ACMIL, DTFD, and HAFED operate on high-resolution feature tensors $U\in\mathbb{R}^{N\times k\times d}$, where $N$ is the number of patches in low-resolution, $k$(16) the number of sub-patches in high-resolution per a patch in low-resolution, and $d$(384) the feature dimension. In contrast, SASHA begins with low-resolution features $Z\in\mathbb{R}^{N\times d}$ and iteratively refines them by zooming into diagnostically relevant patches. The final state representation $S_{T}\in\mathbb{R}^{N\times d}$ can be archived as a compact summary of the WSI, requiring approximately 16 times less storage compared to methods like ACMIL, DTFD, or HAFED, Fig. 3d.
@@ A.4 Computation of WSI compressability, Inference time and ECE
Inference time computation: As shown in Figure 3c, SASHA achieves the lowest inference time compared to models such as ACMIL, DTFD, and HAFED. We report the average inference time per slide in the CAMELYON16 test dataset, inference time includes all steps: pre-processing, feature extraction, and classification.
Inference of all the models ran on an NVIDIA GeForce GTX 1080 Ti GPU. For feature extraction, a batch size of 512 was used.
Expected Calibration Error (ECE): Expected Calibration Error (ECE) is a metric used to quantify how well a neural network’s predicted probabilities align with the true probabilities (or observed frequencies) of its predictions. In simpler terms, it measures how reliable a model’s confidence levels are in its predictions. A perfectly calibrated model would have an ECE of 0, indicating that the model’s confidence scores accurately reflect the true likelihood of its predictions being correct.
To compute ECE, we first divide the predicted confidence scores (i.e., the maximum softmax probabilities) into $M=10$ equally spaced bins. For each bin, we calculate two quantities:
  • Accuracy: the proportion of correctly classified samples in the bin.
  • Confidence: the average of the predicted probabilities in the bin.
These metrics are then used to compute the calibration error for each bin as the absolute difference between its accuracy and confidence. The overall ECE is a weighted average of these differences across all bins, as defined in Equation 2. In below equations $M$ is the number of confidence bins, $B_{m}$ denotes the set of indices of samples falling into the $m^{\text{th}}$ bin, $n$ is the total number of samples, $\hat{y}_{i}$ and $y_{i}$ are the predicted and true labels, respectively, and $\hat{p}_{i}$ is the predicted confidence score for sample $i$ (i.e., the maximum softmax probabilities).
$\text{acc}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\mathbf{1}(\hat{y}_{i}=y_{i})$
$\text{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{p}_{i}$
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|$ (2)
HAFED achieves the lowest Expected Calibration Error (ECE) of approximately 0.015, indicating highly calibrated predictions. In contrast, ACMIL and SASHA-0.1 exhibit the highest ECE values, around 0.06. SASHA-0.2 achieves an ECE of approximately 0.03, which is significantly lower than that of ACMIL and DTFD. This suggests that SASHA-0.2 not only matches HAFED in terms of classification accuracy but also performs comparably in terms of calibration.

### v3

@@ A.3 Additional Control Experiments
A.3.1 Multiclass Classification Performance of SASHA
We extended the SASHA framework to evaluate its performance on multiclass classification tasks using the Camelyon+ and BRACS datasets (see Section˜A.1 for a description of the datasets).
For both datasets, the SASHA framework was trained using experimental settings identical to those of the binary classification tasks. The same patch-level sampling strategy and aggregation mechanism were employed to ensure consistency in evaluation. To mitigate class imbalance, we employed a resampling mechanism that oversamples minority classes (e.g., ITC for the Camelyon+ dataset) during training.
The performance comparison (Table˜9) shows that the methods (HAFED, SASHA) perform comparably with other sota methods, even in this multiclass setting. As before, sampling only 20% of the patches (SASHA-0.2) achieves performance that is comparable to methods that sample all of the WSI patches at high resolution (e.g., DTFD, ACMIL).
Method CAMELYON+ BRACs
Accuracy AUC F1 Accuracy AUC F1
100%
[…]
Table 9: Comparison of multiclass classification performance of different methods on the Camelyon+ and BRACS datasets. Reported metrics include weighted-averaged Accuracy, AUC, and F1-score.
A.3.2 Alternative Models under an Observation Budget
We compared the performance of SASHA with competing sota models like ACMIL [zhang2023acmil] and DTFD [zhang2022dtfd] under the same observation budget of either 10% or 20%. That is, we predicted slide labels by having these sota models examine only a limited fraction of patches in each WSI at high resolution, matching the fraction of the WSI sampled by SASHA.
Each model was trained with a combination of low-resolution and high-resolution patch features and tested, similarly, with a combination of high and low resolution patch features in the same proportion as of training. Specifically, for a given slide, if $N$ is the total number of low-resolution patches and each low-resolution patch corresponds to $k$ high-resolution patches, then for that slide, the input comprised of a total of $N_{L}+N_{H}$ patches comprising $N_{L}=\left(N-\left\lceil f\cdot N\right\rceil\right)$ low-resolution and $N_{H}=\left\lceil f\cdot N\right\rceil\cdot k$ high-resolution patches, where $f\in\{0.1,0.2\}$. Because low and high-resolution patches were of the same pixel dimensionality $(256,256,3)$, the pretrained ViT model (Section˜3) was used to extract features from each such patch; this yielded a feature vector $W\in\mathbb{R}^{(N_{L}+N_{H})\times d}$ which was provided as input to each model for label prediction, both during training and evaluation.
@@ A.3.2 Alternative Models under an Observation Budget
[…]
@@
As shown in Table˜10, such a strategy of sparsely sampling patches substantially reduced the performance of ACMIL and DTFD models. In each case, SASHA sampling either 10% (0.1 fraction) or 20% (0.2 fraction) of the patches at high resolution outperformed the sota models with the corresponding, identical observation budget. In other words, competing sota models failed to live up to SASHA’s performance under a comparable observation budget for high-resolution patches.
A.3.3 Alternative Action Policies during Inference
In the current SASHA implementation, the inference strategy follows a deterministic, greedy policy wherein, for a given state, the action with the highest probability is always selected. This means the agent always picks the most likely action as per the learned policy distribution during inference (but not during training, see Section˜A.2.3).
@@ A.3.3 Alternative Action Policies during Inference
Comparison of different sampling strategies is shown in Table˜11. Interestingly, the default deterministic policy of SASHA – a greedy strategy of selecting the action with the highest probability – outperformed all other stochastic policies. This behavior was observed empirically, where the agent consistently selects action with highest probability given current state $S_{t}$, resulting in more informative state transitions from $S_{t}$ to $S_{t+1}$ and improved classification performance, as shown in Table˜11 . In addition, the stochastic policy that had the greatest degree of freedom for sampling (Stochastic - full distribution) outperformed, by and large, the other policies (top-$k$ and top-$p$); this was understandable because the latter policies impose more constraints on the action space for sampling.
@@
A.3.4 Alternative Reward Schedules during Training
The default SASHA model employs a step-wise reward during training, where the agent receives a reward at every time step of the episode associated with the label prediction loss at that time step (see Section˜A.2.3). In contrast, a previous competing model (RLogist [zhao2022rlogist]) employed a terminal reward strategy, providing feedback only at the end of each episode.
We tested this alternative reward schedule employing only a terminal reward at the end of each episode, both during training and inference. As shown in Table˜12, the RL agent trained with reward at each time step significantly outperformed an agent trained using terminal rewards. We propose that this occurs because our reward formulation – which is designed to minimize the classification loss at each time step – encourages the model to sample more tumor-relevant patches during inference. Confirming this prediction SASHA’s default reward schedule sampled, on average, more patches with a high attention (relevance) score, as compared to the terminal reward schedule (Table˜12).
Reward Schedule Accuracy AUC F1
Top-k attention score
patches $\cap$ RL actions
Avg attention
score of RL actions
@@ A.3.4 Alternative Reward Schedules during Training
Table 12: Comparison of a terminal reward schedule (row 1) with SASHA’s default schedule of reward at each time step (row 2), during training (see text for details). Column 5 tells the average patches per slide that are selected by RL agent are in top-k according to the attention score given by HAFED model. Column 6 tells the average attention score of the selected patches by RL agent. (see Fig.˜3f-g for details). Values denote the performance of the SASHA-0.2 model on the CAMELYON16 dataset. Bold: the best performing reward schedule.
A.3.5 Analysis of RL Agent Behavior
Figure 4: Evaluation of model performance on tumor slides from the test set, categorized into four groups. The first three columns correspond to correctly classified tumor slides (TS) by both SASHA and HAFED across increasing tumor fraction (TF) ranges, while the fourth column represents TS misclassified by SASHA but correctly predicted by HAFED. The x-axis denotes the visited patch fraction, binned with a width of 0.005. For each slide, we compute the average entropy, average tumor class prediction probability, and maximum cumulative tumor patch hit ratio (Eq.˜1) within each bin. These metrics are then averaged across all slides, with solid lines indicating the mean and shaded regions representing the standard deviation.
@@ A.3.5 Analysis of RL Agent Behavior
First, we asked: how quickly does the RL agent visit tumor patches, and how does the model’s confidence evolve over time steps in an episode? We first performed this analysis for slides that were correctly classified by SASHA, separately for 3 classes of WSIs with varying tumor fractions (TF, Fig.˜4, columns 1-3). For this, we defined a cumulative tumor patch Hit Ratio using
[…]
@@
For both variants of the model (SASHA-0.1 and SASHA-0.2), on the slides with very sparse tumor regions (TF<0.1), the classifier’s output favoring the tumor class reaches a high value ($\sim$0.8-1.0) even before it vists 5% of all patches at high-resolution (Fig.˜4, column 1, green trace). In addition, its predictive entropy drops progressively as more patches are sampled (Fig.˜4, column 1, green trace). Surprisingly, even though the peak of the classifier output is reached earlier for SASHA-0.1 as compared to the SASHA-0.2 model, a longer observation budget produces higher accuracies, as well as better calibrated predictions, as shown in the main text Table˜1 and Fig.˜3-h.
As the budget increases from 10% to 20%, the tumor hit ratio improves correspondingly. With a 10% budget, SASHA tends to misclassify slides with small tumor regions (see F1 score in Table˜1), highlighting the importance of increasing the sampling budget. At 20%, SASHA correctly classifies nearly all slides that were correctly predicted by HAFED, demonstrating improved robustness across varying tumor fractions.
As the fraction of tumor patches increases (0.1<TF<0.6 and TF>0.6), the model output favoring the tumor class reaches peak upon visiting progressively miniscule fractions of high-resolution patches (<0.025% or 0.01%, Fig.˜4, columns 2-3, green traces). The predictive entropy also goes to near zero at these fractions indicating confident model predictions with very few samples.
For Fig.˜4, column 2—where the tumor fraction (TF) lies in the range $0.1\leq\text{TF}<0.6$—the tumor hit ratio increases slightly as the budget increases. However, the improvement is minimal because most tumor patches are already captured within the initial 10% of patches selected by the RL agent. This suggests that for WSIs with moderate tumor presence, visiting just 10% of the patches is often sufficient.
@@
We illustrate this with an example slide Fig.˜5. For this slide, the RL agent successfully identifies tumor patches, with the first tumor patch being visited at a fraction of $0.0085$ and the last at $0.03$. For the entire cohort of such slides, tumor patch visits ranged from 0.0010.0990.001-0.099 fraction for SASHA-0.1 and 0.0090.0340.009-0.034 fraction for SASHA-0.2 (Fig.˜4, column 4, orange trace). In contrast, when patches are randomly sampled for the same slide, the first tumor patch is typically sampled after visiting $0.141\pm 0.121$ fraction of the patches. This demonstrates that the RL agent is able to localize tumor regions, typically with an order of magnitude fewer steps than random sampling.
@@
Method Patch FE(HR) FE(LR) Infer TSU RL Total
DTFD 1.46 116.75 N/A 0.0230 N/A N/A 116.77
ACMIL 1.46 116.75 N/A 0.0140 N/A N/A 116.76
HAFED 1.46 116.75 N/A 0.0010 N/A N/A 116.75
SASHA-0.1 1.46 8.60 4.74 0.0007 2.06 0.105 15.50
SASHA-0.2 1.46 18.23 4.95 0.0007 4.09 0.219 27.49
Table 13: Comparison of inference time for different methods, including the time involved in each computational step. “Patch” denotes the time for dividing the slide into patches, FE denotes Feature Extraction time, with HR representing High Resolution and LR representing Low Resolution. “Infer” refers to the inference time for a single model pass. TSU corresponds to the time required for state update, and RL indicates the time needed for agent policy and value updates. The “Total” time (last column) corresponds to the values shown in Fig.˜3c, which does not include the fixed “Patch” time overhead. For SASHA, the value in “Infer.” corresponds specifically to the time for one forward pass through the HAFED model.
A.4 Computation of WSI Compressibility, Inference time, and Calibration Error
Whole Slide Images (WSIs) are typically acquired at gigapixel resolution. For laboratories or hospitals maintaining long-term archives, this poses a significant storage challenge. However, we propose that, for accurate tumor diagnosis, it is not necessary to store WSIs as raw images. By storing them as informative feature representations, a considerable reduction in storage requirements can be achieved.
To quantify the data reduction achieved during WSI processing, we define the WSI compressibility factor as the ratio between the original image size at full magnification and the size of the final feature representation used for classification. We define the raw image size as $W\times H\times 3$, representing the width, height, and RGB color channels, respectively.
Models such as ACMIL, DTFD, and HAFED operate on high-resolution feature tensors $U\in\mathbb{R}^{N\times k\times d}$, where $N$ is the number of low-resolution patches, $k$ (16) is the number of sub-patches in high-resolution per low-resolution patch, and $d$ (384) is the feature dimension. In contrast, SASHA operates on an initial low-resolution feature set $Z\in\mathbb{R}^{N\times d}$ and iteratively refines it by zooming into diagnostically relevant patches. The final state representation $S_{T}\in\mathbb{R}^{N\times d}$ can be archived as a compact summary of the WSI, requiring approximately 16 times less storage compared to methods like ACMIL, DTFD, or HAFED, Fig.˜3d.
@@ A.4 Computation of WSI Compressibility, Inference time, and Calibration Error
Inference time computation: As shown in Figure˜3c, SASHA, with its selective sampling strategy, achieves the lowest inference time compared to models such as ACMIL, DTFD, and HAFED, which sample all the patches at high resolution. We report the average inference time per slide for the CAMELYON16 test dataset. The inference time reported in Fig.˜3c includes all steps: pre-processing, feature extraction, and classification. Table˜13 shows the inference time involved in each step. The column “Patch” refers to the time required to divide the WSI at low resolution into non-overlapping $\times256$ patches, while also removing background content such as regions with only the glass slide or insufficient biological tissue. This step is a constant overhead across all methods and is not included in the Total time in Table˜13. Feature extraction of high-resolution patches reflects a major inference time bottleneck. Because SASHA visits only a small fraction (10-20%) of patches at high resolution, this yields considerable savings over other approaches that visit all (100%) of patches at high resolution, like DTFD, ACMIL, HAFED. In addition to feature extraction, the TSU is a secondary bottleneck in SASHA because at each timestep, the cosine similarity with all other non-masked patches must be computed, albeit at low resolution. Improving the speed of this step is scope for future work.
Expected Calibration Error (ECE): Expected Calibration Error (ECE) is a metric used to quantify how well a neural network’s predicted probabilities align with the true probabilities (or observed frequencies) of its predictions [guo2017calibration]. It measures how reliable a model’s confidence levels are in its predictions. A perfectly calibrated model would have an ECE of 0, indicating that the model’s confidence scores accurately reflect the true likelihood of its predictions being correct.
To compute ECE, we first divide the predicted confidence scores (i.e., the maximum softmax probabilities) into $M=10$ equally spaced bins. For each bin, we calculate two quantities: i) Accuracy: the proportion of correctly classified samples in the bin; and ii) Confidence: the average of the predicted probabilities in the bin. These metrics are then used to compute the calibration error as the absolute difference between its accuracy and confidence for each bin, summed across bins [guo2017calibration].
HAFED achieves the lowest Expected Calibration Error (ECE) of approximately 0.015, indicating highly calibrated predictions. In contrast, ACMIL and SASHA-0.1 exhibit the highest ECE values, around 0.06. SASHA-0.2 achieves an intermediate ECE of around 0.03, which is still lower than that of ACMIL and DTFD. This suggests that SASHA-0.2 not only matches HAFED in terms of classification accuracy but also performs better than other sota methods in terms of calibrated predictions.